# Rethinking Approximate Gaussian Inference in Classification

### v2

@@ 3.2 Intermediate Beta Matching
Figure 2: Approximating Gaussian pushforwards $\Phi_{*}\mathrm{f}(x)$ and $\rho_{*}\mathrm{f}(x)$ through normCDF and sigmoid respectively by Beta distributions $\mathrm{p}(\mathrm{f}(x))$ with moment matching. With normCDF, we can match exact moments, whereas with sigmoid, we match approximate moments. Moreover, when $\mathrm{f}(x)$ is the standard Normal $\mathcal{N}(0,1)$ (on the left in blue), the approximation for the pushforward through normCDF is exact (Section D.1).
When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Eq. 20) further:
@@
where $\mathcal{B}((0,1))$ is the space of Beta distributions on $(0,1)$. Like the output Dirichlets, we obtain the intermediate Beta distributions by moment matching (Appendix D). Figure 2 indicates that such Beta approximations tend to be of high quality. In Section D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions. The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations. Analysing this theoretically presents an interesting direction for future research.
@@ 4 Training
The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix H). Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [37, 19]. The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section 6).
@@ 5 Computational Gains
The runtime cost of the MC predictive approximation is $\mathcal{O}(S\cdot C+C^{3})$, and the memory cost is $\mathcal{O}(S\cdot C+C^{2})$, where $S$ is the number of samples. The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such a covariance matrix before sampling. For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian means an approximately 7% overhead on the forward pass. In embedded systems with much smaller models, this overhead is even larger. In contrast, the runtime and memory cost of calculating our predictives is $\mathcal{O}(C)$, which is negligible in comparison.
Indeed, the approximation Eq. 12 uses only the mean and the diagonal of the logit covariance to calculate the predictive. That is, the logit quantities required to compute a predictive scale as $\mathcal{O}(C)$ instead of $\mathcal{O}(C^{2})$. This can provide further opportunities for computational savings in the methods used to obtain the Gaussian distributions.
In HET, for example, the logit covariance is trained with both a diagonal term and a low-rank term. Our framework thus allows one to drop the low-rank term (Section I.2).
@@ 6 Experiments
    Do we have to sacrifice performance for sample-free predictives? (Section 6.1)
@@
    What are the effects of changing the learning objective? (Section 6.2)
To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives. In the main paper, we report results on Softmax ($\varphi=\exp$, trained with vanilla CE), NormCDF ($\varphi=\Phi$) and Sigmoid ($\varphi=\rho$) models. In Section K.1, we investigate the performance of Softmax models with closed-form predictives.
@@
We consider Heteroscedastic Classifiers (HET) [5], Spectral-Normalized Gaussian Processes (SNGP) [20], and last-layer Laplace-Approximated Networks [8] as backbones supplying logit-space Gaussian distributions. The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [9], CIFAR-10, and CIFAR-100 [17] on five metrics aligning with practical needs from uncertainty estimates [26]. Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples. For ImageNet, we treat ImageNet-C [13] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [14] and Mean Field [21] approximations, as well as MC sampling, in Section 6.2.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones. For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [37] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following Mucsányi et al. [27]. For Vision Transformer [11] experiments, refer to Section K.2. On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [33] in Weights & Biases [3].
The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10. For a complete overview of CIFAR-10 and CIFAR-100 results, refer to Appendix J and Section K.3, respectively.
@@ 6.1 Quality of Sample-Free Predictives
In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions. To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [29]—NormCDF and Sigmoid Laplace (see Fig. 4(a))—and evaluate their per-predictive performance. As Table 1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper. Section K.4 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet. Table 2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
@@ 6.2 Effects of Changing the Learning Objective
The previous section shows that for models trained with Sigmoid or NormCDF, our closed-form predictives always perform on par with MC sampling while being more efficient. However, our proposed objective (see Eq. 29) changes the training dynamics and, subsequently, the performance of models. In this section, we showcase the performance of methods equipped with our closed-form predictives and learning objectives against Softmax models. For the latter, we use the best-performing predictive and estimator (see Appendices E and F for an overview of predictive approximations and estimators). We employ our methods with the closed-form predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
We first evaluate how calibrated the models are using the test NLL [12] and Expected Calibration Error (ECE) [29] metrics. Fig. 3(b) shows that on CIFAR-10, the score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods. On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig. 3(a)).
@@
The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy. Fig. 4(a) shows that, on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
Next, we turn to the correctness prediction task: whether the models can predict the correctness of their own predictions. We consider correctness estimators $t(x):=\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictive mean $\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[\mathbf{P}\right]$. Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness, defined as $v(x)=[\hat{y}(x)=y]\in\{0,1\}$ with $\hat{y}(x):=\arg\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in\{1,\dots,C\}$. We then measure the binary log probability score of $\tilde{t}(x)$, given by $v(x)\log t(x)+(1-v(x))\log(1-t(x))$. Fig. 4(b) shows that our closed-form predictives outperform all Softmax predictives across all methods.
@@
Fig. 5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives. These results are in line with the findings of Wightman et al. [38], who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function. Section K.5 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture thereof. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator. As OOD inputs, we consider corrupted ImageNet-C [13] samples. Fig. 5(b) shows that our closed-form predictives outperform Softmax across all methods. Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section 5). See Section K.6 for the other severity levels.
@@ Appendix A Gaussian Integral Derivations
In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Eq. 8) and normCDF (Eq. 13), as well as the approximations for pushforwards through sigmoid (Eq. 15) and softmax.
@@ A.1 Gaussian Exp Integral
[…]
@@ A.2 Gaussian NormCDF Integral
Here, we derive the classical normCDF Gaussian integration formula [30, Eq. 10,010.8].
@@ A.3 Probit Approximation for Gaussian Sigmoid Integral
Taylor expanding $\rho$ to first order about 00,
@@
as $y\to 0$. Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$. So using Eq. 31 we derive the probit approximation [34, 23]
@@ A.4 Mean Field Approximation for Gaussian Softmax Integral
For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [21] is obtained as follows
@@
(1) is the mean field approximation, and (3) uses the probit approximation Eq. 33. [21] provides two other variants of this approximation with other choices of approximation (2).
@@ Appendix B Comparison with the Multinomial Probit Model
In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [6].
@@
In the case $C=2$, the multinomial probit model Eq. 37 outputs closed-form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Eq. 37 and Eq. 38:
@@ Appendix C Theoretical Analyses
In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations. This complements the empirical analyses, for instance in the synthetic experiment (Fig. 1) or in [7].
@@
which is well defined if $\bm{p}$, $\bm{q}$ lie in the interior of the simplex ($p_{i},q_{i}\neq 0,1$ for all $i$). So, for a predictive approximation, $\bm{\hat{p}}$, we would like to analyse $\operatorname{D_{KL}}(\bm{p},\bm{\hat{p}})$, where $\bm{p}:=\mathbb{E}_{\bm{P}\sim\bm{a}_{*}\mathcal{N}(\bm{\mu},\bm{\Sigma})}[\bm{P}]$ is the true predictive, $\bm{\mu}$ and $\bm{\Sigma}$ are some logit space mean and covariance and $\bm{a}$ is an output activation (e.g. $\bm{a}=\bm{n}\circ\bm{\varphi}$).

### v3

@@ 4.2 Intermediate Beta Matching
Figure 2: Approximating Gaussian pushforwards $\Phi_{*}\mathrm{f}(x)$ and $\rho_{*}\mathrm{f}(x)$ through normCDF and sigmoid respectively by Beta distributions $\mathrm{p}(\mathrm{f}(x))$ with moment matching. With normCDF, we can match exact moments, whereas with sigmoid, we match approximate moments. Moreover, when $\mathrm{f}(x)$ is the standard Normal $\mathcal{N}(0,1)$ (on the left in blue), the approximation for the pushforward through normCDF is exact (Section˜D.1).
When $\varphi=\Phi\text{ or }\rho$, noting that their range is $(0,1)$, we can decompose the approximate distributional model (Eq.˜20) further:
@@
where $\mathcal{B}((0,1))$ is the space of Beta distributions on $(0,1)$. Like the output Dirichlets, we obtain the intermediate Beta distributions by moment matching (Appendix˜D). Figure˜2 indicates that such Beta approximations tend to be of high quality. In Section˜D.1, we give an information-geometric motivation for this, which shows that the choice of activation $\Phi$ is ideal for approximating the Gaussian pushforwards with Beta distributions. The quality of such Beta approximations is encouraging for that of the output Dirichlet approximations. Analysing this theoretically presents an interesting direction for future research.
@@ 5 Training
The BCE loss is a valid choice of loss as it is a (negative) strictly proper scoring rule (see Appendix˜I). Moreover, it has been observed to provide advantageous training dynamics compared to the CE loss [rw2019timm, li_bce_2025]. The choice of $\varphi=\Phi$ or $\rho$ not only fits our framework for closed-form predictives, but allows one to leverage such advantageous training dynamics (see experimental results in Section˜7).
@@
In contrast to the previously mentioned methods, HET directly trains not only for the mean $\bm{f}$ but also for the covariance matrix $\bm{\Sigma}$ of $\mathrm{f}$ through the predictives. Thus, in this case, we use the BCE with the pushforward means $\bar{Q}_{c}:=\mathbb{E}_{Q_{c}\sim\varphi_{*}\mathrm{f}_{c}(x)}[Q_{c}]$:
$\mathcal{L}((x_{n},c_{n})_{n=1}^{N})=-\sum_{n=1}^{N}\sum_{c=1}^{C}\big(\delta_{c_{n},c}\log(\bar{Q}_{c})+(1-\delta_{c_{n},c})\log(1-\bar{Q}_{c})\big).$ (30)
@@ 6 Computational Gains
The runtime cost of the MC predictive approximation is $\mathcal{O}(S\cdot C+C^{3})$, and the memory cost is $\mathcal{O}(S\cdot C+C^{2})$, where $S$ is the number of samples. The $C^{2}$ term in memory comes from the logit covariance matrix; the $C^{3}$ term in runtime comes from the need to Cholesky-decompose such a covariance matrix before sampling. For example, on a ResNet-50, sampling one thousand logit vectors from the logit-space Gaussian incurs approximately 7% overhead on the forward pass: see Appendix˜H for MC timing experiments. In embedded systems with much smaller models, this overhead is even larger. In contrast, the runtime and memory cost of calculating our predictives is $\mathcal{O}(C)$, which is negligible in comparison.
Indeed, the approximation Eq.˜12 uses only the mean and the diagonal of the logit covariance to calculate the predictive. That is, the logit quantities required to compute a predictive scale as $\mathcal{O}(C)$ instead of $\mathcal{O}(C^{2})$. This provides further opportunities for computational savings in the methods used to obtain the Gaussian distributions. E.g., in HET, the logit covariance is trained with both a diagonal term and a low-rank term. Our framework thus allows dropping the low-rank term (Section˜J.2).
@@ 7 Experiments
    Do we have to sacrifice performance for sample-free predictives? (Section˜7.1)
@@
    What are the effects of changing the learning objective? (Section˜7.2)
To answer the first question, we consider fixed (method, activation) pairs and verify that our methods perform on par with the Monte Carlo sampled predictives. In the main paper, we report results on Softmax ($\varphi=\exp$, trained with vanilla CE), NormCDF ($\varphi=\Phi$) and Sigmoid ($\varphi=\rho$) models. In Section˜K.1, we investigate the performance of Softmax models with closed-form predictives.
@@
We consider Heteroscedastic Classifiers (HET) [collier2021correlated], Spectral-Normalised Gaussian Processes (SNGP) [liu2020simple], and last-layer Laplace-Approximated Networks [daxberger2021laplace] as backbones supplying logit-space Gaussian distributions. The resulting 21 (method, activation, predictive) triplets are evaluated on ImageNet-1k [deng2009imagenet], CIFAR-10, and CIFAR-100 [krizhevsky2009learning] on five metrics aligning with practical needs from uncertainty estimates [mucsanyi2023trustworthy]. Further, we test the Out-of-Distribution (OOD) detection capabilities of the models on balanced mixtures of In-Distribution (ID) and OOD samples. For ImageNet, we treat ImageNet-C [hendrycks2019benchmarking] samples with 15 corruption types and 5 severity levels as OOD samples. For CIFAR-10 and CIFAR-100, we use the CIFAR-10C corruptions.
For the second question, we evaluate our closed-form predictives and moment-matched Dirichlets against softmax models with the Laplace Bridge [pmlr-v180-hobbhahn22a] and Mean Field [lu_mean-field_2021] approximations, the linear-time approximation by shekhovtsov_feed-forward_2018, and MC sampling; see Section˜7.2.
To provide a fair comparison, we reimplement each method as simple-to-use wrappers around deterministic backbones. For ImageNet evaluation, we use a ResNet-50 backbone from the timm library [rw2019timm] pretrained with the softmax activation function, and train each (method, activation) pair for 50 ImageNet-1k epochs following mucsányi2024benchmarking. For Vision Transformer [dosovitskiy2020image] experiments, refer to Section˜K.2. On CIFAR-10, we train ResNet-28 models from scratch for 100 epochs. The CIFAR-100 experiments train WideResNet-28-5 models for 200 epochs. We search for ideal hyperparameters with a ten-step Bayesian Optimization scheme [shahriari2015taking] in Weights & Biases [wandb].
The main paper focuses on ImageNet results to highlight the scalability of our framework and only shows proper scoring results for CIFAR-10. For a complete overview of CIFAR-10 and CIFAR-100 results, refer to Section˜K.3 and Section˜K.4, respectively.
@@ 7.1 Quality of Sample-Free Predictives
In this section, we investigate whether there is a price to pay for sample-free predictives and second-order distributions. To this end, we take the two best-performing methods on the Expected Calibration Error (ECE) metric [naeini2015obtaining]—NormCDF and Sigmoid Laplace (see Fig.˜4(a))—and evaluate their per-predictive performance. As Table˜1 shows, the closed-form predictive is always on par with the (unbiased) MC estimates on ImageNet while being cheaper. Section˜K.5 reports the KL divergence of different approximations to the ‘true’ predictive (estimated using 10,000 MC samples) on ImageNet. Table˜2 shows that the moment-matched second-order Dirichlet distributions consistently outperform both MC and sample-free predictives on the OOD detection task, showcasing the practical utility of second-order distributional knowledge.
@@ 7.2 Effects of Changing the Learning Objective
The previous section shows that for models trained with Sigmoid or NormCDF, our closed-form predictives always perform on par with MC sampling while being more efficient. However, our proposed objective (see Eq.˜29) changes the training dynamics and, subsequently, the performance of models. In this section, we showcase the performance of methods equipped with our closed-form predictives and learning objectives against Softmax models. For the latter, we use the best-performing predictive and estimator (see Appendices˜E and F for an overview of predictive approximations and estimators). We employ our methods with the closed-form predictives and second-order Dirichlet distributions to demonstrate their competitive nature while being more efficient than MC sampling.
We first evaluate how calibrated the models are using the test NLL [gneiting2007strictly] and Expected Calibration Error (ECE) [naeini2015obtaining] metrics. Fig.˜3(b) shows that on CIFAR-10, the NLL score of our closed-form predictives (Sigmoid, NormCDF) is consistently better than the corresponding Softmax results for all methods. On the large-scale ImageNet dataset, Sigmoid and NormCDF predictives either outperform or are on par with Softmax (see Fig.˜3(a)). The ECE metric requires the models’ confidence (i.e., maximum class probability) to match their accuracy. Fig.˜4(a) shows that on ImageNet, our closed-form predictives have a clear advantage over Softmax across all methods.
@@
Next, we turn to the correctness prediction task: whether the models can predict the correctness of their own predictions. We consider correctness estimators $t(x):=\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in[0,1]$ for inputs $x\in\mathcal{X}$ derived from the predictive mean $\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[\mathbf{P}\right]$. Framed as a binary prediction task, the goal of these estimators is to predict the probability of the predicted class’ correctness, defined as $v(x)=[\hat{y}(x)=y]\in\{0,1\}$ with $\hat{y}(x):=\arg\max_{1\leq c\leq C}\mathbb{E}_{\mathbf{P}\sim\mathrm{h}(x)}\left[P_{c}\right]\in\{1,\dots,C\}$. We then measure the binary log probability score of $\tilde{t}(x)$, given by $v(x)\log t(x)+(1-v(x))\log(1-t(x))$. Fig.˜4(b) shows that our closed-form predictives outperform all Softmax predictives across all methods.
@@
Fig.˜5(a) evidences this claim on ImageNet: our closed-form predictives either outperform or are on par with Softmax predictives. These results are in line with the findings of wightman_resnet_2021, who also recommend training with a per-class binary cross-entropy loss using the sigmoid activation function. Section˜K.6 shows that the BCE loss already yields performance gains in the model backbones (without the logit-space Gaussians).
Finally, we consider another binary prediction task, where a general uncertainty estimator $u(x)\in\mathbb{R}$ (derived from predictives or second-order Dirichlet distributions) is tasked to separate ID and OOD samples from a balanced mixture thereof. As the uncertainty estimator can take on any real value, we measure the Area Under the Receiver Operating Characteristic curve (AUROC), which quantifies the separability of ID and OOD samples w.r.t. the uncertainty estimator. As OOD inputs, we consider corrupted ImageNet-C [hendrycks2019benchmarking] samples. Fig.˜5(b) shows that our closed-form predictives outperform Softmax across all methods. Importantly, these closed-form predictives and second-order Dirichlet distributions are considerably cheaper to calculate than Softmax MC predictions (see Section˜6). See Section˜K.7 for the other severity levels.
@@ Appendix A Gaussian Integral Derivations
In this appendix, we derive the closed-form formula for the mean of Gaussian pushforwards through exp (Eq.˜8) and normCDF (Eq.˜13), as well as the approximations for pushforwards through sigmoid (Eq.˜15) and softmax.
@@ A.1 Gaussian Exp Integral
$\int_{\mathbb{R}}\exp(y)\;\mathcal{N}(\mu,\sigma^{2})(dy)$ $=\int_{\mathbb{R}}\exp(y)\exp\left(-\frac{(y-\mu)^{2}}{2\sigma^{2}}\right)\;dy$ (31)
$=\int_{\mathbb{R}}\exp\left(\mu+\frac{\sigma^{2}}{2}\right)\exp\left(-\frac{1}{2\sigma^{2}}(y-\mu-\sigma^{2})^{2}\right)\;dy$
@@ A.2 Gaussian NormCDF Integral
Here, we derive the classical normCDF Gaussian integration formula [owen_table_1980, Eq. 10,010.8].
@@ A.3 Probit Approximation for Gaussian Sigmoid Integral
Taylor expanding $\rho$ to first order about 0,
@@
as $y\to 0$. Hence matching $\rho$ and $\Phi$ to first order we get the approximation $\rho(y)\approx\Phi\left(\sqrt{\frac{\pi}{8}}y\right)$. So using Eq.˜32 we derive the probit approximation [spiegelhalter_sequential_1990, mackay_evidence_1992]
@@ A.4 Mean Field Approximation for Gaussian Softmax Integral
For $\bm{\mu}\in\mathbb{R}^{C}$, $\bm{\sigma}^{2}\in\mathbb{R}^{C}_{>0}$ and $\bm{\Sigma}=\operatorname{\mathbf{diag}}(\bm{\sigma}^{2})$, the mean field approximation to the Gaussian softmax integral [lu_mean-field_2021] is obtained as follows
@@
(1) is the mean field approximation, and (3) uses the probit approximation Eq.˜34. [lu_mean-field_2021] provides two other variants of this approximation with other choices of approximation (2).
@@ Appendix B Comparison with the Multinomial Probit Model
In this appendix, we show that a model whose output activation is a composition of an element-wise normCDF activation $\bm{\Phi}$ and a normalisation $\bm{n}$ is distinct from the classical multinomial probit model [daganzo_chapter_1979].
@@
In the case $C=2$, the multinomial probit model Eq.˜38 outputs closed-form probabilities. This allows us to construct an explicit counterexample to the equivalence of the two models Eq.˜38 and Eq.˜39:
@@ Appendix C Theoretical Analyses
In this appendix, we provide (formal and informal) theoretical analyses of the quality of various predictive approximations. This complements the empirical analyses, for instance in the synthetic experiment (Fig.˜1) or in [daunizeau_semi-analytical_2017].
@@
[…]